OpenStudy (anonymous):

Explain, in complete sentences, how you would divide the following expression: . 12rt x^5 x 6rt x Write the simplified answer in radical form.

OpenStudy (anonymous):

\[\sqrt[12]{5} \times \sqrt[6]{x}\]